find the slope and the-intercept of the 5x=6y+36

Answer by nerdybill(7384)

Rearrange:
5x=6y+36
so that it is in "slope-intercept" form:
y = mx + b
.
5x=6y+36
0=6y-5x+36
-6y=-5x+36
6y=5x-36
y=(5/6)5x-6
From inspection then:
slope is 5/6
y-intercept is at (0,-6)